# Linear fractional relative risk aversion

### v1

@@ Introduction
Assume now that $\alpha\neq 0$. Then, we can establish the if part by using (5). Using the differentiation formula for the hypergeometric function $\frac{\partial}{\partial(-\alpha\sigma q)}\,{}_{2}F_{1}\bigl{(}1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}=\frac{1-\frac{1}{\sigma}}{2-\beta}\,{}_{2}F_{1}\bigl{(}2-\frac{1}{\sigma},2;3-\beta;-\alpha\sigma q\bigr{)}$, we have $\frac{\partial}{\partial q}\,{}_{2}F_{1}\bigl{(}1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}=(-\alpha\sigma)\frac{1-\frac{1}{\sigma}}{2-\beta}\,{}_{2}F_{1}\bigl{(}2-\frac{1}{\sigma},2;3-\beta;-\alpha\sigma q\bigr{)}$. [^7] See https://dlmf.nist.gov/15.5.1 (NIST Digital Library of Mathematical Functions). We can thus differentiate (5) with respect to $q$ to obtain
$u^{\prime}(q)$ $=$ $\textstyle\frac{K\sigma}{\sigma-1}[1-\beta+\alpha(\sigma-1)q](1+\alpha\sigma q)^{-1+\beta-\frac{1}{\sigma}}q^{-\beta}\left[1+\frac{\beta-\frac{1}{\sigma}}{1-\beta}\,{}_{2}F_{1}\left(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\right)\right]$ (6)
@@
The contiguous relationship of the hypergeometric functions implies: [^8] See https://dlmf.nist.gov/15.5.19 (NIST Digital Library of Mathematical Functions).
@@
Plugging this into (6), we have the marginal utility function as follows.
@@
$u(q)=\textstyle c_{1}+c_{2}(-\alpha\sigma q)^{1-\beta}{}_{2}F_{1}\bigl{(}\frac{1}{\sigma}-\beta,1-\beta;2-\beta;-\alpha\sigma q\bigr{)}.$ (7)
@@
Applying this transformation to (7) yields
$u(q)=\textstyle c_{1}+c_{2}(-\alpha\sigma q)^{1-\beta}(1+\alpha\sigma q)^{1-\frac{1}{\sigma}+\beta}{}_{2}F_{1}\bigl{(}2-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}.$ (8)
@@
Solving it for $(1+\alpha\sigma q){}_{2}F_{1}\bigl{(}2-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}$, plugging the resulting expression into (8), and noting that $_{2}F_{1}\bigl{(}1-\frac{1}{\sigma},0;2-\beta;-\alpha\sigma q\bigr{)}=1$ yield:
@@
Hence, we have three types of quantity restrictions for the utility function in (5), i.e., $u^{\prime}(q)>0$, $u^{\prime\prime}(q)<0$, and $\lvert-\alpha\sigma q\rvert<1$, where the last condition is required from the definition of $_{2}F_{1}\left(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\right)$ in (5). We first summarize the first two restrictions in Assumption 1 (see Table 1). We then show in Proposition 1 that we can use analytic continuation to extend the utility function (5) to all quantities satisfying the restrictions in Table 1, so that the third restriction $\lvert-\alpha\sigma q\rvert<1$ is immaterial.
Assumption 1
  $\alpha\geq 0$ $\alpha<0$
$\alpha\sigma\geq 0$   $q\in[0,\infty)$ $q\in[0,-\frac{\beta}{\alpha})$
Notes: The CRRA case with $\beta=\frac{1}{\sigma}$ requires $\sigma>1$ for $\beta\in(0,1)$, which implies that either the top-left cell or the bottom-right cell is applicable. In the latter case, $q\in[0,-\frac{1}{\alpha\sigma})\cap[0,-\frac{\beta}{\alpha})$ must be replaced with $q\in[0,\infty)$ (see footnote 12).
Proposition 1 (Integral representation)
Proof.
@@
$_{2}F_{1}\bigl{(}1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}$ in (5) is defined for $\lvert-\alpha\sigma q\rvert<1$. Thus, there are two cases depending on the sign of $\alpha\sigma$. [^13] Strictly speaking, there is another case with $\alpha\sigma=0$, which implies $\alpha=0$ by the assumption that $\sigma\neq 0$ in Theorem 1 and Proposition 1. Since $_{2}F_{1}\bigl{(}1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}=1$ for $\alpha=0$, we do not have to replace $_{2}F_{1}\bigl{(}1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}$ in (5) with the integral in (11). However, we can let $\alpha\to 0$ in (11) to obtain $u(q)=\frac{Kq^{1-\beta}}{1-\beta}+C$ for all $q\in[0,\infty)$, which is consistent with Table 1 and equivalent to (5) evaluated at $\alpha=0$. We will return to this case in Corollary 1. First, when $\alpha\sigma<0$, $\lvert-\alpha\sigma q\rvert<1$ implies $0\leq q<-\frac{1}{\alpha\sigma}$, which is not restrictive given Table 1. Second, when $\alpha\sigma>0$, $\lvert-\alpha\sigma q\rvert<1$ implies $0\leq q<\frac{1}{\alpha\sigma}$, which appears to be restrictive given Table 1. However, since $2-\beta>1>0$ holds in Theorem 1, we can extend $_{2}F_{1}\bigl{(}1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}$ in (5) to $q\geq\frac{1}{\alpha\sigma}$ by analytic continuation (see, e.g., (9.1.6) in Lebedev, 1965, where $\lvert{\rm arg}(1-z)\rvert=\lvert{\rm arg}(1+\alpha\sigma q)\rvert=0$ since $1+\alpha\sigma q>0$). Thus, in both cases, we can use the integral representation $_{2}F_{1}\bigl{(}1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}=\frac{\Gamma(2-\beta)}{\Gamma(1)\Gamma(1-\beta)}\int_{0}^{1}(1-t)^{-\beta}(1+t\alpha\sigma q)^{-(1-\frac{1}{\sigma})}{\rm d}t$, where $\Gamma(\cdot)$ is the gamma function. Using the recurrence relation $\Gamma(2-\beta)=(1-\beta)\Gamma(1-\beta)$ and $\Gamma(1)=1$, the foregoing expression reduces to $_{2}F_{1}\bigl{(}1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr{)}=(1-\beta)\int_{0}^{1}(1-t)^{-\beta}(1+t\alpha\sigma q)^{-(1-\frac{1}{\sigma})}{\rm d}t$. Hence, for all $q$ satisfying the quantity restrictions in Table 1, we can rewrite (5) as (11). $\square$
Theorem 1 does not include the two corner cases $\beta=0$ and $\beta=1$. The next theorem extends Theorem 1 to the corner cases and shows that the utility function that has LFRRA nests the HARA and CREMR utility functions as special cases. [^14] The HARA utility (e.g., Merton, 1971) is defined either as satisfying that the inverse of the absolute risk aversion is linear in $q$, i.e., $\frac{u^{\prime}(q)}{u^{\prime\prime}(q)}=\frac{K(1+\alpha\sigma q)^{-\frac{1}{\sigma}}}{\alpha K(1+\alpha\sigma q)^{-\frac{1}{\sigma}-1}}=\frac{1+\alpha\sigma q}{\alpha}$ or that the prudence $\frac{u^{\prime\prime\prime}(q)}{u^{\prime\prime}(q)}$ is proportional to the absolute risk aversion $\frac{u^{\prime\prime}(q)}{u^{\prime}(q)}$, i.e., $\frac{u^{\prime\prime\prime}(q)u^{\prime}(q)}{[u^{\prime\prime}(q)]^{2}}=1+\sigma$.
Theorem 2 (Limit cases)
Proof.
@@
To establish (12) and (13), we first rewrite (5). Recall that (5) is derived from a linear combination of the two generally linearly independent solutions $_{2}F_{1}\bigl{(}\frac{1-\sigma}{\sigma},0;\beta;z\bigr{)}$ and $z^{1-\beta}{}_{2}F_{1}\bigl{(}1+\frac{1-\sigma}{\sigma}-\beta,1-\beta;2-\beta;z\bigr{)}$. However, when $\beta$ goes to 1 from below, these solutions are no longer linearly independent. Thus, to obtain both (12) and (13) from a single expression, we use the following transformation of the hypergeometric function for $z=-\alpha\sigma q\notin(1,\infty)$: [^15] See https://functions.wolfram.com/07.23.17.0079.01 for that transformation. Observe that the condition $z=-\alpha\sigma q\notin(1,\infty)$ always holds for all quantities satisfying the restrictions in Table 1.
@@
Plugging this expression into the second term of (2) we obtain
$u(q)$ $=$ $\textstyle\frac{K\sigma(1+\alpha\sigma q)^{\beta-\frac{1}{\sigma}}q^{1-\beta}}{\sigma-1}-\frac{K\sigma}{\sigma-1}\frac{\Gamma(1-\beta)\Gamma(\frac{1}{\sigma})}{\Gamma(-\beta+\frac{1}{\sigma})}\bigl{(}\frac{1}{\alpha\sigma}\bigr{)}^{1-\beta}$ (15)
@@
$=$ $\textstyle\frac{K\sigma(1+\alpha\sigma q)^{\beta-\frac{1}{\sigma}}q^{1-\beta}}{\sigma-1}\bigl{[}1+\frac{1-\beta\sigma}{\alpha\sigma q}{}_{2}F_{1}\bigl{(}1,\beta;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr{)}\bigr{]}-\frac{K\sigma}{\sigma-1}\frac{\Gamma(1-\beta)\Gamma(\frac{1}{\sigma})}{\Gamma(-\beta+\frac{1}{\sigma})}\bigl{(}\frac{1}{\alpha\sigma}\bigr{)}^{1-\beta}+C,\qquad$
where $_{2}F_{1}\bigl{(}1,\beta;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr{)}=\sum_{n=0}^{\infty}\frac{(1)_{n}(\beta)_{n}}{(1+\frac{1}{\sigma})_{n}}\frac{(-\frac{1}{\alpha\sigma q})^{n}}{n!}$, with $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$, denotes the hypergeometric function.
Using (15), we now establish (12) and (13). First, when $\beta$ goes to 0 from above, we can disregard the restriction $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$ because we have $\lim_{\beta\to 0^{+}}{}_{2}F_{1}\bigl{(}1,\beta;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr{)}=1$, regardless of the value of $\frac{1}{\alpha\sigma q}$. Hence, letting $\beta$ in (15) go to 0 from above and by noting that $\lim_{\beta\to 0^{+}}{}_{2}F_{1}\bigl{(}1,\beta;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr{)}=1$ and $\Gamma(1)=1$, we obtain (12) as follows:
@@
Second, to establish (13), let $C=\frac{K\sigma}{\sigma-1}\frac{\Gamma(1-\beta)\Gamma(\frac{1}{\sigma})}{\Gamma(-\beta+\frac{1}{\sigma})}\bigl{(}\frac{1}{\alpha\sigma}\bigr{)}^{1-\beta}+\widetilde{C}$ in (15). We then have
@@
$u^{\prime}(q)=K(1+\alpha\sigma q)^{-\frac{1}{\sigma}}\quad\mbox{and}\quad u^{\prime}(q)=K(1+\alpha\sigma q)^{1-\frac{1}{\sigma}}q^{-1},$ (16)
For the marginal utility to be strictly decreasing, we impose $\alpha>0$ in the former and $1+\alpha q>0$ in the latter. [^16] These conditions correspond to $u^{\prime\prime}(q)=-K(\alpha q+\beta)(1+\alpha\sigma q)^{-1+\beta-\frac{1}{\sigma}}q^{-1-\beta}<0$ that we assume for the case $\beta\in(0,1)$ in Theorem 1. Indeed, letting $\beta\to 0^{+}$ and $\beta\to 1^{-}$ in this inequality, we obtain $\alpha>0$ and $1+\alpha q>0$. If $\alpha>0$, these conditions are satisfied for all $q\geq 0$. If $\alpha<0$, the latter condition implies $0\leq q<-\frac{1}{\alpha}$ since $q\geq 0$ must hold. Thus, we can extend the ranges of quantities for $\beta\in(0,1)$ in Table 1 to include the limit cases with $\beta\to 0^{+}$ and $\beta\to 1^{-}$.
Hence, as in Theorem 1, we have three types of quantity restrictions for the utility function (13), i.e., $u^{\prime}(q)>0$, $u^{\prime\prime}(q)<0$, and $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$, where the last condition is required from the definition of $_{2}F_{1}\bigl{(}1,1;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr{)}$ in terms of the power series. We have shown that the first two restrictions are consistent with Table 1. The last restriction, $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$, requires $\alpha\sigma>0$. [^17] We can show that $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$ requires $\alpha\sigma>0$ as follows. If $\alpha\sigma<0$ were to hold, then $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$ would imply $\frac{1}{\alpha\sigma q}<1$, so that $1<-\alpha\sigma q$, thus violating our assumption $1+\alpha\sigma q>0$. The case with $\alpha\sigma=0$ is excluded by the assumptions in Theorem 2. We will elaborate on this case in Corollary 1. It can also be used to show that $1+\frac{1}{\sigma}>0$, which allows us to exclude the possibility that $1+\frac{1}{\sigma}$ can be a non-positive integer, in which case $_{2}F_{1}\bigl{(}1,1;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr{)}$ is not well defined. [^18] If $\sigma>0$, we readily obtain $1+\frac{1}{\sigma}>0$. If $\sigma<0$, we have $\alpha<0$ since $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$ requires $\alpha\sigma>0$ as shown in footnote 17. In that case, the restriction $1+\alpha q>0$ implies $q<-\frac{1}{\alpha}$. Since $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$ implies $\frac{1}{\alpha\sigma}<q$, it follows that $\frac{1}{\alpha\sigma}<q<-\frac{1}{\alpha}$, which is a non-empty interval only if $\frac{1}{\alpha\sigma}<-\frac{1}{\alpha}$. This, together with $\alpha<0$, implies $\frac{1}{\sigma}>-1$. Hence, we obtain $1+\frac{1}{\sigma}>0$. Recall that $\sigma=0$ is excluded in Theorem 2. We can summarize these restrictions in Assumption 2.
Assumption 2
Proposition 2 (Integral representation)
Proof.

### v2

@@ Appendix A.1 Proof of Theorem 1
Assume now that $\alpha\neq 0$. Then, we can establish the if part by using (5). Using the differentiation formula for the Gauss hypergeometric function $\frac{\partial}{\partial(-\alpha\sigma q)}\,{}_{2}F_{1}\bigl(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr)=\frac{1-\frac{1}{\sigma}}{2-\beta}\,{}_{2}F_{1}\bigl(2-\frac{1}{\sigma},2;3-\beta;-\alpha\sigma q\bigr)$, we have $\frac{\partial}{\partial q}\,{}_{2}F_{1}\bigl(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr)=(-\alpha\sigma)\frac{1-\frac{1}{\sigma}}{2-\beta}\,{}_{2}F_{1}\bigl(2-\frac{1}{\sigma},2;3-\beta;-\alpha\sigma q\bigr)$. [^26] See https://dlmf.nist.gov/15.5.1 (NIST Digital Library of Mathematical Functions). We can thus differentiate (5) with respect to $q$ to obtain
$u^{\prime}(q)$ $=$ $\textstyle\frac{K\sigma}{\sigma-1}[1-\beta+\alpha(\sigma-1)q](1+\alpha\sigma q)^{-1+\beta-\frac{1}{\sigma}}q^{-\beta}\left[1+\frac{\beta-\frac{1}{\sigma}}{1-\beta}\,{}_{2}F_{1}\left(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\right)\right]$ (A.1)
@@
The contiguous relationship of the Gauss hypergeometric functions implies: [^27] See https://dlmf.nist.gov/15.5.19 (NIST Digital Library of Mathematical Functions).
@@
Plugging this into (A.1), we have the marginal utility function as follows:
@@
$u(q)=\textstyle c_{1}+c_{2}(-\alpha\sigma q)^{1-\beta}{}_{2}F_{1}\bigl(\frac{1}{\sigma}-\beta,1-\beta;2-\beta;-\alpha\sigma q\bigr).$ (A.2)
@@
Applying this transformation to (A.2) yields
$u(q)=\textstyle c_{1}+c_{2}(-\alpha\sigma q)^{1-\beta}(1+\alpha\sigma q)^{1-\frac{1}{\sigma}+\beta}{}_{2}F_{1}\bigl(2-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr).$ (A.3)
@@
Solving it for $(1+\alpha\sigma q){}_{2}F_{1}\bigl(2-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr)$, plugging the resulting expression into (A.3), and noting that $_{2}F_{1}\bigl(1-\frac{1}{\sigma},0;2-\beta;-\alpha\sigma q\bigr)=1$ yield:
@@
Appendix A.2 Proof of Proposition 1
@@
The Gauss hypergeometric function $_{2}F_{1}\bigl(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr)$ in (5) is defined for $\lvert-\alpha\sigma q\rvert<1$. Thus, there are two cases depending on the sign of $\alpha\sigma$. [^31] Strictly speaking, there is another case with $\alpha\sigma=0$, which implies $\alpha=0$ by the assumption that $\sigma\neq 0$ in Theorem 1 and Proposition 1. Since $_{2}F_{1}\bigl(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr)=1$ for $\alpha=0$, we do not have to replace $_{2}F_{1}\bigl(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr)$ in (5) with the integral in (8). However, we can let $\alpha\to 0$ in (8) to obtain $u(q)=\frac{Kq^{1-\beta}}{1-\beta}+C$ for all $q\in[0,\infty)$, which is consistent with Table 1 and equivalent to (5) evaluated at $\alpha=0$. We will return to this case in Corollary 1. First, when $\alpha\sigma<0$, $\lvert-\alpha\sigma q\rvert<1$ implies $0\leq q<-\frac{1}{\alpha\sigma}$, which is not restrictive given Table 1. Second, when $\alpha\sigma>0$, $\lvert-\alpha\sigma q\rvert<1$ implies $0\leq q<\frac{1}{\alpha\sigma}$, which appears to be restrictive given Table 1. However, since $2-\beta>1>0$ holds in Theorem 1, we can extend $_{2}F_{1}\bigl(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr)$ in (5) to $q\geq\frac{1}{\alpha\sigma}$ by analytic continuation (see, e.g., (9.1.6) in Lebedev, 1965, where $\lvert{\rm arg}(1-z)\rvert=\lvert{\rm arg}(1+\alpha\sigma q)\rvert=0$ since $1+\alpha\sigma q>0$). Thus, in both cases, we can use the integral representation $_{2}F_{1}\bigl(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr)=\frac{\Gamma(2-\beta)}{\Gamma(1)\Gamma(1-\beta)}\int_{0}^{1}(1-t)^{-\beta}(1+t\alpha\sigma q)^{-(1-\frac{1}{\sigma})}{\rm d}t$, where $\Gamma(\cdot)$ is the gamma function. Using the recurrence relation $\Gamma(2-\beta)=(1-\beta)\Gamma(1-\beta)$ and $\Gamma(1)=1$, the foregoing expression reduces to $_{2}F_{1}\bigl(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\bigr)=(1-\beta)\int_{0}^{1}(1-t)^{-\beta}(1+t\alpha\sigma q)^{-(1-\frac{1}{\sigma})}{\rm d}t$. Hence, for all $q$ satisfying the quantity restrictions in Table 1, we can rewrite (5) as (8). $\square$
Appendix A.3 Proof of Theorem 2
@@
To establish (9) and (10), we first rewrite (5). Recall that (5) is derived from a linear combination of the two generally linearly independent solutions $_{2}F_{1}\bigl(\frac{1-\sigma}{\sigma},0;\beta;z\bigr)$ and $z^{1-\beta}{}_{2}F_{1}\bigl(1+\frac{1-\sigma}{\sigma}-\beta,1-\beta;2-\beta;z\bigr)$. However, when $\beta$ goes to 1 from below, these solutions are no longer linearly independent. Thus, to obtain both (9) and (10) from a single expression, we use the following transformation of the Gauss hypergeometric function for $z=-\alpha\sigma q\notin(1,\infty)$: [^32] See https://functions.wolfram.com/07.23.17.0079.01 for that transformation. Observe that the condition $z=-\alpha\sigma q\notin(1,\infty)$ always holds for all quantities satisfying the restrictions in Table 1.
@@
Plugging this expression into the second term of (Appendix A.3) we obtain
$u(q)$ $=$ $\textstyle\frac{K\sigma(1+\alpha\sigma q)^{\beta-\frac{1}{\sigma}}q^{1-\beta}}{\sigma-1}-\frac{K\sigma}{\sigma-1}\frac{\Gamma(1-\beta)\Gamma(\frac{1}{\sigma})}{\Gamma(-\beta+\frac{1}{\sigma})}\bigl(\frac{1}{\alpha\sigma}\bigr)^{1-\beta}$ (A.5)
@@
$=$ $\textstyle\frac{K\sigma(1+\alpha\sigma q)^{\beta-\frac{1}{\sigma}}q^{1-\beta}}{\sigma-1}\bigl[1+\frac{1-\beta\sigma}{\alpha\sigma q}{}_{2}F_{1}\bigl(1,\beta;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr)\bigr]-\frac{K\sigma}{\sigma-1}\frac{\Gamma(1-\beta)\Gamma(\frac{1}{\sigma})}{\Gamma(-\beta+\frac{1}{\sigma})}\bigl(\frac{1}{\alpha\sigma}\bigr)^{1-\beta}+C,\qquad\ \$
where $_{2}F_{1}\bigl(1,\beta;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr)=\sum_{n=0}^{\infty}\frac{(1)_{n}(\beta)_{n}}{(1+\frac{1}{\sigma})_{n}}\frac{(-\frac{1}{\alpha\sigma q})^{n}}{n!}$, with $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$, denotes the Gauss hypergeometric function.
Using (A.5), we now establish (9) and (10). First, when $\beta$ goes to 0 from above, we can disregard the restriction $\lvert-\frac{1}{\alpha\sigma q}\rvert<1$ because we have $\lim_{\beta\to 0^{+}}{}_{2}F_{1}\bigl(1,\beta;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr)=1$, regardless of the value of $\frac{1}{\alpha\sigma q}$. Hence, letting $\beta$ in (A.5) go to 0 from above and by noting that $\lim_{\beta\to 0^{+}}{}_{2}F_{1}\bigl(1,\beta;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr)=1$ and $\Gamma(1)=1$, we obtain (9) as follows:
@@
Second, to establish (10), let $C=\frac{K\sigma}{\sigma-1}\frac{\Gamma(1-\beta)\Gamma(\frac{1}{\sigma})}{\Gamma(-\beta+\frac{1}{\sigma})}\bigl(\frac{1}{\alpha\sigma}\bigr)^{1-\beta}+\widetilde{C}$ in (A.5). We then have
@@
Appendix A.4 Proof of Proposition 2